Determine the force in each member of the Fink roof truss shown using Method of Joints. State whether each member is in tension or compression. Forces A= 5 kN, B= 10kN, D= 10 kN , F = 10 kN ,and G = 5kN2.25 m 2.25 m KN A 3 m3 m3 m
Answer #1

Draw the FBD of the truss. TO KN 2.25m TO KN OKN 5KN SKN . Efx=0: Ax=0 AX 3m ć 3m E 3m 1 Since the loading is symmetric w Ay10KN Joint | wy? Foo 2.25 7+ S FBC FAB=36-93 KN + E Fx = 0 + 2.25 S Fan + 2-25 3.4620 2.462 (36.33) - 0.6 Foc +0.913 8 FA - -EFxzo Fce - 33.75 + 3/5 (9.374) + 3% (9.374) 20 Fee = 22.50 kn) (tention) Due to Symmetry! . FoE - Fco =) FDE = 9.374 KN (ren